# A Fractional Variational Approach to Spectral Filtering Using the Fourier Transform

### v1

@@ 2.2 Fractional Filter
where $\alpha$ is a real number satisfying $n-1\leq\alpha\leq n$, with $n\in\mathbb{N}$, and $t>a>0$ is the independent variable, in which $a$ is the lower bound.
@@
  • (e)
    The fractional derivative is an ill-posed operator; small errors in the input can lead to large deviations in the output. The ill-conditioning depends on the order of the fractional derivative: the higher the order, the more severe the ill-conditioning.
Fractional derivatives have well-defined representations in the frequency domain. Specifically, the Fourier transform of the Riemann–Liouville derivative of order $\alpha$ is given by
@@ 2.2.1 Two dimension version
The fractional Laplace operator $(-\Delta)^{\alpha/2}f$ is also known as the Riesz derivative, $^{R}D^{\alpha}f$, with $0<\alpha\leq 2$ [29, 30]. The most direct and operational definition of the fractional Laplacian is in the frequency domain, through the Fourier transform, in which case we have
@@
is the Euclidean norm of the frequency vector. For $\alpha=2$, we have $(-\Delta)^{\alpha/2}=-\Delta$. This definition arises naturally in various contexts. [^3] When the domain $\Omega$ is bounded and sufficiently regular, the eigenfunctions $\{\phi_{k}\}$ and the eigenvalues $\{\lambda_{k}\}$ of the Dirichlet Laplacian satisfy $\lambda_{k}>0$ and $\begin{cases}-\Delta\phi_{k}=\lambda_{k}\phi_{k},&\text{in }\Omega,\\[4.0pt] \phi_{k}=0,&\text{on }\partial\Omega,\end{cases}$ where $\{\phi_{k}\}$ form an orthonormal set. This spectral decomposition allows one to define the fractional powers of the Laplacian by $(-\Delta)^{s}u:=\sum_{k}\lambda_{k}^{s}\,u_{k}\,\phi_{k}.$ In the specific case $\Omega=[0,1]^{2}$ the eigenfunctions are $\phi_{m,n}(x,y)=\sin(m\pi x)\sin(n\pi y),\qquad m,n\in\mathbb{N},$ with corresponding eigenvalues $\lambda_{m,n}=\pi^{2}(m^{2}+n^{2}).$
The fractional Laplacian is preferred over the “fractional gradient” because the Laplacian is a scalar operator (acting from $\mathbb{R}^{n}$ to $\mathbb{R}$), while the gradient is a vector. When the Laplacian is raised to a fractional power, the scalar structure is preserved, which greatly facilitates both analysis and application.

### v2

@@ 2.2 Fractional Filter
where $\alpha$ is a real number satisfying $n-1\leq\alpha\leq n$, with $n\in\mathbb{N}$, and $t>a>$-$\infty$ is the independent variable, in which $a$ is the lower bound. In this work, we adopt $a=$-$\infty$ and the Liouville fractional derivative, for which the Fourier transform is well defined.
@@
Fractional derivatives have well-defined representations in the frequency domain. Specifically, the Fourier transform of the Liouville derivative of order $\alpha$ is given by
@@ 2.2.1 Two dimension version
The most direct and operational definition of the fractional Laplacian is in the frequency domain, through the Fourier transform, in which case we have
@@
is the Euclidean norm of the frequency vector. For $\alpha=2$, we have $(-\Delta)^{\alpha/2}=-\Delta$. This definition arises naturally in various contexts. [^3] When the domain $\Omega$ is bounded and sufficiently regular, the eigenfunctions $\{\phi_{k}\}$ and the eigenvalues $\{\lambda_{k}\}$ of the Dirichlet Laplacian satisfy $\lambda_{k}>0$ and $\begin{cases}-\Delta\phi_{k}=\lambda_{k}\phi_{k},&\text{in }\Omega,\\[4.0pt] \phi_{k}=0,&\text{on }\partial\Omega,\end{cases}$ where $\{\phi_{k}\}$ form an orthonormal set. This spectral decomposition allows one to define the fractional powers of the Laplacian by $(-\Delta)^{s}u:=\sum_{k}\lambda_{k}^{s}\,u_{k}\,\phi_{k}.$ In the specific case $\Omega=[0,1]^{2}$ the eigenfunctions are $\phi_{m,n}(x,y)=\sin(m\pi x)\sin(n\pi y),\qquad m,n\in\mathbb{N},$ with corresponding eigenvalues $\lambda_{m,n}=\pi^{2}(m^{2}+n^{2}).$ However, this is neither the only nor necessarily the physically most appropriate way to define a fractional Laplacian [33]. The drawback of taking this as a starting point is that the Fourier transform is no longer available for bounded domains, although it remains applicable in this setting for zero boundary conditions.
The fractional Laplacian is preferred over the “fractional gradient” because the Laplacian is a scalar operator (acting from $\mathbb{R}^{n}$ to $\mathbb{R}$), while the gradient is a vector. When the Laplacian is raised to a fractional power, its scalar structure is preserved, which greatly facilitates both analysis and applications. Moreover, the fractional Laplacian admits multiple characterizations on different domains, and there is currently no consensus in the literature regarding the most appropriate definition of the fractional Laplacian [34, 35].
@@ 3.1.2 Shannon Entropy
Another parameter that can be used to assess the image is the Tsallis entropy, which can more effectively capture long-range correlations, non-Gaussian behavior, and more complex textures. In the reference [36], the Tsallis entropy was computed from the two-dimensional histogram constructed using both the gray-level values of the pixels and their local average gray values. Reference [36] demonstrates that, in terms of the quality of the thresholded images, Tsallis entropy yields better results than the Shannon method.